y"+(x-1)y'+y=0


1
Expert's answer
2021-02-24T06:56:45-0500

"y''+(x-1)+y=0"

Applying Power series solution.

Let

"y=\sum_{n=0}^\infty a_nx^n"

So

"y'=\sum_{n=1}^\infty na_nx^{n-1}\\
y''=\sum_{n=2}^\infty n(n-1)a_nx^{n-2}"

Substitute these values for "y,y'" and "y''" into the DE.

"\sum_{n=2}^\infty n(n-1)a_nx^{n-2}+(x-1)\sum_{n=1}^\infty na_nx^{n-1}+\sum_{n=0}^\infty a_nx^n=0\\
\sum_{n=2}^\infty n(n-1)a_nx^{n-2}+x\sum_{n=1}^\infty na_nx^{n-1}-\sum_{n=1}^\infty na_nx^{n-1}+\sum_{n=0}^\infty a_nx^n=0\\
\sum_{n=2}^\infty n(n-1)a_nx^{n-2}+\sum_{n=1}^\infty na_nx^n-\sum_{n=1}^\infty na_nx^{n-1}+\sum_{n=0}^\infty a_nx^n=0\\
\sum_{n=0}^\infty (n+1)(n+2)a_{n+2}x^n+\sum_{n=0}^\infty na_nx^n-\sum_{n=0}^\infty (n+1)a_{n+1}x^{n+1}+\sum_{n=0}^\infty a_nx^n=0\\"

Now, equating the coefficients of "x^n" to 0

"(n+1)(n+2)a_{n+2}+na_n-(n+1)a_{n+1}+a_n=0\\
(n+1)(n+2)a_{n+2}=(n+1)a_{n+1}-(n+1)a_n\\
a_{n+2}=\frac{a_{n+1}-a_n}{n+2}"

Now putting "n=0,1,2,\cdots" . We have

"a_2=-\frac{1}{2}(a_0-a_1)\\
a_3=\frac{1}{3}(a_2-a_1)=-\frac{1}{6}(a_0+a_1)\\
a_4=\frac{1}{4}(a_3-a_2)=\frac{1}{12}(a_0-2a_1)\\
\cdots"

Recall that ;

"y(x)=\sum_{n=0}^\infty a_nx^n=a_0+a_1x+a_2x^2+a_3x^3+a_4x^4+\cdots"

We will finally have y(x) in terms of "a_0,a_1."

"y(x)=a_0+a_1x-\frac{1}{2}(a_0-a_1)x^2-\frac{1}{6}(a_0+a_1)x^3+\frac{1}{12}(a_0-2a_1)x^4+\cdots\\
y(x)=a_0(1-\frac{1}{2}x^2-\frac{1}{6}x^3+\frac{1}{12}x^4+\cdots)+a_1(x+\frac{1}{2}x^2-\frac{1}{6}x^3-\frac{1}{6}x^4+\cdots)."